OpenStudy (kwilliam98):

Rachel deposited $6449.82 into a savings account with an interest rate of 6.9% compounded annually. About how long will it take for the account to be worth $8000?

OpenStudy (kwilliam98):

.

OpenStudy (shadelily):

This is a geometric sequence. The formula for that would be \[Un = U _{1} \times r ^{n-1} \ge 8000\] So what you need to do is sub in those values U1 (initial value) = 6449.82 r (interest) = 1.069 Now just sub in and solve for n. \[Un = 6449.82 \times 1.069^{n-1} \ge 8000\] Now simply solve it as an inequality. I hope it helps...